# DeepDetect: Learning All-in-One Dense Keypoints

### v2

@@ Abstract
Keypoint detection is the foundation of many computer vision tasks, including image registration, structure-from-motion, 3D reconstruction, visual odometry, and SLAM. Traditional detectors (SIFT, SURF, ORB, BRISK, etc.) and learning-based methods (SuperPoint, R2D2, LF-Net, D2-Net, etc.) have shown strong performance yet suffer from key limitations: sensitivity to photometric changes, low keypoint density and repeatability, limited adaptability to challenging scenes, and lack of semantic understanding, often failing to prioritize visually important regions. We present DeepDetect, an intelligent, all-in-one, dense keypoint detector that unifies the strengths of classical detectors using deep learning. Firstly, we create ground-truth masks by fusing outputs of 7 keypoint and 2 edge detectors, extracting diverse visual cues from corners and blobs to prominent edges and textures in the images. Afterwards, a lightweight and efficient model: “ESPNet”, is trained using these masks as labels, enabling DeepDetect to focus semantically on images while producing highly dense keypoints, that are adaptable to diverse and visually degraded conditions. Evaluations on the Oxford Affine Covariant Regions dataset demonstrate that DeepDetect surpasses other detectors in keypoint density, repeatability, and the number of correct matches, achieving maximum values of 0.5143 (average keypoint density), 0.9582 (average repeatability), and 59,003 (correct matches).
@@ I Introduction
Keypoint detection is the backbone of many computer vision applications including image matching, panorama stitching, object tracking, 3D reconstruction, visual odometry, and visual SLAM [1]. Keypoints (often referred to as features or feature points) are the distinctive points in an image that can also be identified to an extent under various transformations or degradations by the detectors. They are usually corners, blobs, ridges, and textured patterns that stand out from their surroundings [1]. Selection of keypoint detector is a critical decision in computer vision as it determines the stability, repeatability, map density, and overall performance of the downstream application [2, 3, 4, 5, 6]. A variety of detectors have been proposed over the years, ranging from traditional algorithms (SIFT, SURF, Harris Corner Detector, ORB, BRISK, and AGAST) to deep learning based approaches (SuperPoint, R2D2, LF-Net, and D2-Net). After keypoint detection, corresponding feature descriptors are computed by encoding the local pixel neighborhood around each keypoint into a distinctive vector that enhances stability across image variations [1]. A good descriptor is more distinctive and invariant to diverse image transformations or degradations [7, 8, 9, 10].
@@
Classical detectors are invariant to different image transformations (scale, rotation, affine, blur) but sensitive to extreme variations (intense darkness, extreme photometric/geometric changes, severe smoke or fog) [8]. With standard settings, they detect sparse keypoints that inadequately cover texture-rich and low-contrast regions [8]. They also lack semantic understanding to prioritize important regions in the images. Learning-based detectors (SuperPoint, R2D2, LF-Net, D2-Net) improve keypoint generalization and stability but their qualitative results show similar shortcomings [15, 16, 17, 18]. They detect sparse keypoints and lack semantic awareness, potentially missing important regions in cluttered or low-visibility scenes [15, 16, 17, 18]. These issues highlight the need for an intelligent and dense keypoint detector that is robust to complex image transformations and focuses on important regions.
@@
  • We propose a multi-detector, multi-edge fusion pipeline that generates diverse and rich keypoint supervision masks, which serve as labels for training DeepDetect.
  • We use a lightweight and efficient model: “ESPNet”, to ensure that DeepDetect has a small memory footprint:  1.82 MB, making it well-suited for deployment on edge devices such as mobile phones.
  • We demonstrate that DeepDetect outperforms traditional keypoint detectors both qualitatively and quantitatively by significant margins. On the Oxford dataset [11], DeepDetect achieves the highest performance, with an average repeatability of 0.9582, average keypoint density of 0.5143, and 59,003 correct matches, outperforming the traditional detectors even with their extremely low thresholds.
@@ II Related Work
Classical feature detectors have laid the foundation of modern computer vision. However, they rely on handcrafted criteria such as gradient magnitude, intensity variation, or corner response thresholds. Although computationally efficient, they often fail with standard thresholds under extreme image transformations, lacking adaptability to changing visibility conditions [8]. Learning based methods aim to learn keypoint detection, description (or sometimes both) but they typically produce sparse keypoints [15, 16, 17, 18] which are not suitable for computer vision applications that require dense keypoints such as dense 3D reconstruction, AR/VR, and SLAM.
@@ II-A Classical Detectors and Descriptors
II-B Learning based Detectors and Descriptors
SuperPoint [15] is a well-known learning based detector and descriptor trained via synthetic homographies, offering high repeatability and descriptor quality across varying conditions. R2D2 [16] learns both repeatability and descriptor reliability via deep networks, filtering out unstable keypoints for robust matching. LF-Net [17] jointly learns detection and description in a differentiable pipeline using image pairs and geometric constraints. D2-Net [18] uses a CNN to jointly detect and describe dense keypoints in a single forward pass, particularly effective for challenging photometric and geometric changes. Although having some strengths, these algorithms lack semantic awareness about the scenes and do not provide dense keypoints that are especially required in mapping applications such as dense 3D reconstruction, AR/VR, and visual SLAM.
@@ III DeepDetect – Methodology
III-A Data Preparation
Data preparation is the first stage towards the development of DeepDetect as shown in Figure 3. We selected 41000 images from the MS-COCO [27] and NewTsukuba [28] datasets, applying extreme reductions in brightness and contrast to 25% of them. Afterwards, they were split into training and validation sets of 33000 and 8000 images, respectively. The strengths of the 7 keypoint detectors and 2 edge detectors are then unified by creating a combined binary mask of their outputs for each image. SIFT detects blob-like structures, while the other 6 keypoint detectors identify corners or ridges. Canny highlights fine details, whereas Sobel captures thicker, high-intensity edges in the images. Binary masks are generated with pixels belonging to the detected keypoints and edges marked as 1 and all other pixels as 0. Extremely low detection thresholds are used for the low visibility scenes, whereas for normal scenes moderate thresholds are used. This mitigates irrelevant keypoints from the normal scenes while maximizes their detection in the challenging scenes. For each image, its keypoint and edge masks are merged to create a rich, multi-cue supervision mask that serves as its label during training. Given an image $I\in\mathbb{R}^{H\times W\times 3}$, set of keypoint detectors D = {SIFT, ORB, BRISK, FAST, AGAST, Harris, Shi-Tomasi} and set of edge detectors E = {Canny, Sobel}, the final combined mask M(I), shown in Figure 4(d), is obtained as:
@@ III-A Data Preparation
where $\vee$ denotes the pixel wise logical OR operation, $M_{D}(I)$ shows the fusion mask obtained from the keypoint detectors as in Figure 4(b) and $M_{E}(I)$ shows the fusion mask obtained from the edge detectors as in Figure 4(c). The construction of the final mask can also be expressed as:
$M(I)=\begin{cases}1,&\text{if }\exists\,p\in D\cup E\text{ such that }M_{p}(I)=1\\ 0,&\text{otherwise}\end{cases}\vskip 5.69046pt$ (2)
III-B ESPNet based Training
Efficient Spatial Pyramid Network (ESPNet) [29] is a light weight model designed for highly efficient pixel-wise prediction tasks such as “semantic segmentation” on resource-constrained devices. It has encoder-decoder architecture: the encoder captures contextual information through successive convolution and down-sampling layers while the decoder progressively up-samples the feature maps to restore the spatial resolution. ESPNet is based on ESP modules which factorize the standard convolutions into point-wise $(1\times 1)$ projections followed by spatial pyramids of dilated convolutions [29]. These modules also have residual connections that assist in smooth feature propagation through the model. ESPNet works on a lower dimensional space due to the point-wise projections, reducing its memory footprint and computational cost. Compared to heavy image segmentation models like U-Net [30] and PSPNet [31], it is upto $180\times$ smaller, $22\times$ faster, and an excellent choice for low complexity tasks [29].
In the training stage (see Figure 3), ESPNet is trained on the prepared dataset for $100$ epochs with a learning rate of $0.001$ and batch size of $64$, using cosine annealing scheduler for smoother convergence. We used Python, OpenCV, and PyTorch [32] for coding. The input image size is kept consistent at $3\times 480\times 480$ for the model and its predicted output mask is re-sized to match the spatial dimensions of the original image. Given a target mask $M(I)=y\in\{0,1\}^{H\times W}$ and model’s raw output $z\in\mathbb{R}^{H\times W}$, the BCE loss $L(y,z)$ is described as:
@@ III-C DeepDetect — Dense Keypoint Detection
DeepDetect is not only learned to reproduce the geometric diversity of the classical keypoint and edge detectors but also able to capture semantically important regions, leading to dense and foreground focused keypoints. To perform image matching, keypoints are first detected in the image pair using the trained ESPNet. Afterwards, we use SIFT descriptor to compute the corresponding descriptors, because of its highly distinctive description nature. The computed descriptor sets for the two images are then matched using NNDR based matching. RANSAC can be applied to reject the outliers and filter the correct matches. Qualitative results of DeepDetect are shown in Figure 6. For the comparative analysis presented in Table II and Figure 6, we used the ground-truth Homography matrices [11] (not RANSAC) to obtain the number of correct matches.
@@ 66 Correct Matches
790 Correct Matches
16581 Correct Matches
2715 Correct Matches
59003 Correct Matches
502 Correct Matches
4756 Correct Matches
40204 Correct Matches
Figure 6: Qualitative comparison of DeepDetect with SIFT under default and extremely low [8] thresholds. The yellow lines show correct correspondences between the images. DeepDetect provides substantially higher correct matches, without requiring manual tuning (additional results provided in the appendix).
@@
Results show that DeepDetect achieves state-of-the-art performance. On the Oxford dataset [11], it gives an average repeatability score of 0.9582, surpassing other keypoint detectors by a significant margin. The results for average keypoint density are presented in Table I, where DeepDetect achieves the highest average density of 0.5143. Table I also reports average foreground-to-keypoint ratio on $18$ custom-selected images with ground-truth binary masks. DeepDetect achieves good ratios, indicating that its keypoints are not only dense but also well-concentrated in semantically important regions in the images. With default thresholds, the classical detectors yield higher F-KP ratios but significantly lower keypoint densities. Conversely, reducing thresholds to increase density decreases the F-KP ratio, highlighting the presence of noisy keypoints. For instance, FAST and AGAST with low thresholds, achieve high densities with low F-KP ratios, showing their inability to concentrate on salient regions. DeepDetect produces considerably higher densities while maintaining strong foreground focus. Visual comparisons in Figure 2 clearly show that DeepDetect focuses on semantically meaningful regions and provides very high keypoint density. Table II and Figure 6 show that DeepDetect provides substantially higher correct matches and consistently outperforms classical detectors. The strength of DeepDetect lies in its ability to combine the geometric robustness and diversity of keypoint and edge detectors, and adaptability to extreme image variations gained through the use of supervision masks generated with different thresholds for normal and challenging scenes. DeepDetect demonstrates stronger generalization across degraded visual conditions due to its intelligent nature.

### v3

@@ Abstract
Keypoint detection is the foundation of many computer vision tasks, including image registration, structure-from-motion, 3D reconstruction, visual odometry, and SLAM. Traditional detectors (SIFT, ORB, BRISK, FAST, etc.) and learning-based methods (SuperPoint, R2D2, QuadNet, LIFT, etc.) have shown strong performance gains yet suffer from key limitations: sensitivity to photometric changes, low keypoint density and repeatability, limited adaptability to challenging scenes, and lack of semantic understanding, often failing to prioritize visually important regions. We present DeepDetect, an intelligent, all-in-one, dense detector that unifies the strengths of classical detectors using deep learning. Firstly, we create ground-truth masks by fusing outputs of 7 keypoint and 2 edge detectors, extracting diverse visual cues from corners and blobs to prominent edges and textures in the images. Afterwards, a lightweight and efficient model: “ESPNet”, is trained using fused masks as labels, enabling DeepDetect to focus semantically on images while producing highly dense keypoints, that are adaptable to diverse and visually degraded conditions. Evaluations on Oxford, HPatches, and Middlebury datasets demonstrate that DeepDetect surpasses other detectors achieving maximum values of 0.5143 (average keypoint density), 0.9582 (average repeatability), 338,118 (correct matches), and 842,045 (voxels in stereo 3D reconstruction).
@@ I Introduction
Keypoint detection is the backbone of many computer vision applications including image matching, panorama stitching, object tracking, 3D reconstruction, visual odometry, and visual SLAM [23]. Keypoints (often referred to as features or feature points) are the distinctive points in an image that can also be identified to an extent under various transformations or degradations by the detectors. They are usually corners, blobs, ridges, and textured patterns that stand out from their surroundings [23]. Selection of keypoint detector is a critical decision in computer vision as it determines the stability, repeatability, map density, and overall performance of the downstream application [7, 4, 12, 18, 10]. A variety of detectors have been proposed over the years, ranging from traditional algorithms (SIFT, SURF, Harris Corner Detector, ORB, BRISK, and AGAST) to deep learning based approaches (SuperPoint, R2D2, D2-Net, Quad-Net, and LIFT). After keypoint detection, corresponding feature descriptors are computed by encoding the local pixel neighborhood around each keypoint into a distinctive vector that enhances stability across image variations [23]. A good descriptor is more distinctive and invariant to diverse image transformations or degradations [34, 33, 14, 16].
@@
Classical detectors are invariant to different image transformations (scale, rotation, affine, blur) but sensitive to extreme variations (intense darkness, extreme photometric/geometric changes, severe smoke or fog) [33]. With standard settings, they detect sparse keypoints that inadequately cover texture-rich and low-contrast regions [33]. They also lack semantic understanding to prioritize important regions in the images. Learning-based detectors SuperPoint, R2D2, D2-Net, Quad-Net, and LIFT) improve keypoint generalization and stability but their qualitative results show similar shortcomings [8, 26, 9, 30, 35]. They detect sparse keypoints and lack semantic awareness, potentially missing important regions in cluttered or low-visibility scenes [8, 26, 9, 30, 35]. This highlights the need for an intelligent and dense keypoint detector that is robust to complex image transformations and focuses on important regions.
@@
  • We name this novel concept of keypoint-edge fusion as DeepDetect, that generates rich and diverse supervision masks, which later on serve as labels for training.
  • We use a lightweight, efficient model: “ESPNet”, to ensure that DeepDetect has a small memory footprint:  1.82 MB and  0.41 M parameters, making it well-suited for deployment on edge devices such as mobile phones.
  • We demonstrate that DeepDetect outperforms other detectors both qualitatively and quantitatively by significant margins. DeepDetect provides highest performance across three different datasets, with an average repeatability of 0.9582 (Oxford), average keypoint density of 0.5143 (Oxford), 338,118 number of correct matches (HPatches), and as high as 842,045 voxels in 3D reconstruction (Middlebury), outperforming the classical detectors even when extremely low thresholds are used.
@@ II Related Work
Classical feature detectors have laid the foundation of modern computer vision. However, they rely on handcrafted criteria such as gradient magnitude, intensity variation, or corner response thresholds. Although computationally efficient, they often fail with standard thresholds under extreme image transformations, lacking adaptability to changing visibility conditions [33]. Learning based methods aim to learn keypoint detection, description (or sometimes both) but they typically produce sparse keypoints [8, 26, 9, 30, 35]. which are not suitable for computer vision applications that require dense keypoints such as dense 3D reconstruction, AR/VR, and dense Simultaneous Localization and Mapping (SLAM).
@@ II-A Classical Detectors and Descriptors
II-B Learning based Detection and Description
SuperPoint [8] is a well-known learning based detector and descriptor trained via synthetic homographies, offering high repeatability and descriptor quality across varying conditions. R2D2 [26] learns both repeatability and descriptor reliability via deep learning, filtering out unstable keypoints for robust matching. D2-Net [9] uses a CNN to jointly detect and describe dense keypoints in a single forward pass, particularly effective for challenging photometric and geometric changes. QuadNet is a learning-based detector that uses a quadruplet CNN to learn repeatable and discriminative keypoints without relying on handcrafted heuristics [30]. LIFT [35] learns detection, orientation estimation, and description using three CNNs. Although having some strengths, these algorithms lack semantic awareness about the scenes and do not provide dense keypoints that are especially required in mapping applications such as dense 3D reconstruction, AR/VR, and visual SLAM.
@@ III DeepDetect – Methodology
III-A Data Preparation – Keypoint-Edge Fusion
Data preparation is the first stage towards the development of DeepDetect as shown in Figure 3. We selected 41000 images from the MS-COCO [19] and NewTsukuba [25] datasets, applying extreme reductions in brightness and contrast to 25% of them. Afterwards, they were split into training and validation sets of 33000 and 8000 images, respectively. The strengths of the 7 keypoint detectors and 2 edge detectors are then unified by creating a combined binary mask of their outputs for each image. SIFT detects blob-like structures, while the other 6 keypoint detectors identify corners or ridges. Canny highlights fine details, whereas Sobel captures thicker, high-intensity edges in the images. Binary masks are generated with pixels belonging to the detected keypoints-edges marked as 1 and all other pixels as 0. Extremely low detection thresholds are used for the low visibility scenes, whereas moderate thresholds are used for the normal scenes. This mitigates irrelevant keypoints from the normal scenes while maximizing detection in challenging scenes. For each image, its keypoint and edge masks are merged to create a rich, multi-cue supervision mask that serves as its label during training. Given an image $I\in\mathbb{R}^{H\times W\times 3}$, set of keypoint detectors D = {SIFT, ORB, BRISK, FAST, AGAST, Harris, Shi-Tomasi} and set of edge detectors E = {Canny, Sobel}, the final combined mask M(I), shown in Figure 4(d), is obtained as:
@@ III-A Data Preparation – Keypoint-Edge Fusion
where $\vee$ denotes the pixel wise logical OR operation, $M_{D}(I)$ shows the fusion mask obtained from the keypoint detectors as in Figure 4(b) and $M_{E}(I)$ shows the fusion mask obtained from the edge detectors as in Figure 4(c).
III-B Learning All-in-One Keypoints – Training Stage
ESPNet [22] is a light weight model designed for highly efficient pixel-wise prediction tasks such as “semantic segmentation” on resource-constrained devices. It has encoder-decoder architecture: the encoder captures contextual information through successive convolution and down-sampling layers while the decoder progressively up-samples the feature maps to restore the spatial resolution. ESPNet is based on ESP modules which factorize the standard convolutions into point-wise $(1\times 1)$ projections followed by spatial pyramids of dilated convolutions [22]. ESPNet works on a lower dimensional space due to the point-wise projections, reducing its memory footprint and computational cost. Compared to heavy image segmentation models like U-Net [27] and PSPNet [36], it is upto $180\times$ smaller, $22\times$ faster, and an excellent choice for low complexity tasks [22].
In the training stage (Figure 3), ESPNet is trained on the prepared dataset for $100$ epochs with a learning rate of $0.001$ and batch size of $64$, using cosine annealing scheduler for smooth convergence. We used Python, OpenCV, and PyTorch [24] for coding. The input image size is kept consistent at $3\times 480\times 480$ for the model and its predicted output mask is re-sized to match the spatial dimensions of the original image. Given a target mask $M(I)=y\in\{0,1\}^{H\times W}$ and model’s raw output $z\in\mathbb{R}^{H\times W}$, the BCE loss $L(y,z)$ is described as:
@@ III-C DeepDetect — Dense Keypoint Detection
DeepDetect is not only learned to reproduce the geometric diversity of the classical keypoint and edge detectors but also able to capture semantically important regions, leading to dense and foreground focused keypoints. To perform image matching, keypoints are first detected in the image pair using the trained ESPNet. Afterwards, we use SIFT descriptor to compute the corresponding descriptor vectors, because of its highly distinctive nature. The descriptor sets for the two images are then matched using NNDR based matching. RANSAC can be applied to reject the outliers and filter the correct matches. Qualitative results of DeepDetect are shown in Figure 6. For the comparative analysis presented in Table II and Figure 6, we used the ground-truth Homography matrices [13] (not RANSAC) to obtain the number of correct matches.
@@ IV Experiments & Results
Trees-1 and Trees-6 [13]
50 Correct Matches
368 Correct Matches
2,202 Correct Matches
@@ 39 Correct Matches
1,020 Correct Matches
6,517 Correct Matches
7,822 Correct Matches
@@ 66 Correct Matches
16,581 Correct Matches
2,715 Correct Matches
47,447 Correct Matches
59,003 Correct Matches
@@
4,756 Correct Matches
40,204 Correct Matches
9,368 Correct Matches
Figure 6: Qualitative comparison of DeepDetect with SIFT under default and extremely low [33] thresholds. Yellow lines show correct correspondences between the images. DeepDetect provides dense and substantially higher correct matches, without requiring manual tuning (due to its intelligent nature).
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
Figure 7: Qualitative comparison of DeepDetect, LIFT [35], and SuperPoint [8] on HPatches [2]. Green lines show the correct matches between the image pairs: (astronautis_1_4, boutique_1_6, ajuntament_1_5, home_1_3, dogman_1_6, nescafe_1_5, crownnight_1_2, busstop_1_6). DeepDetect provides dense and substantially higher correct matches across all scenes, achieving over 0.24 million and 0.33 million matches for the final two image pairs, respectively.
@@
$\text{Foreground-to-Keypoint Ratio}=\frac{N_{F}}{N_{T}}$ (6)
Cones [32]
Dolls [31]
Moebius [31]
Rocks [31]
Cloth [31]
550 Voxels
3,788 Voxels
2,527 Voxels
5,230 Voxels
5,270 Voxels
14,128 Voxels
39,557 Voxels
47,016 Voxels
80,529 Voxels
67,973 Voxels
49,979 Voxels
540,780 Voxels
363,703 Voxels
709,425 Voxels
842,045 Voxels
Figure 8: Comparison of DeepDetect with SIFT for 3D reconstruction (using only two images per scene) on Middlebury datasets [32, 31]. Model used here is trained only on normally illuminated images to avoid low-visibility bias, boosting DeepDetect’s performance for scene geometry. Each row (left to right), shows the left camera image of scene, followed by the 3D reconstructed point clouds using SIFT (default thresholds), SIFT (extremely low thresholds) [33], and DeepDetect. DeepDetect produces highly dense 3D reconstructions, outperforming SIFT even when extremely low thresholds are used.
TABLE III: Comparison of DeepDetect’s repeatability with four well-known learning based detection and description algorithms (MagicPoint [8], SuperPoint [8], QuadNet [30], and R2D2 [26]) on HPatches dataset [2] and Oxford dataset [13].
HPatches Dataset MagicPoint SuperPoint DeepDetect
57 Illumination Scenes 0.575 0.652 0.896
59 Viewpoint Scenes 0.322 0.503 0.901
Oxford Dataset QuadNet R2D2 DeepDetect
Graf 0.25 0.47 0.89
Wall 0.46 0.71 0.98
Bark 0.16 0.47 0.91
Boat 0.29 0.57 0.99
Leuven 0.77 0.77 0.99
Bikes 0.57 0.76 0.94
Trees 0.50 0.60 0.95
UBC 0.68 0.68 0.90
Results show that DeepDetect achieves state-of-the-art performance. On Oxford dataset [13], it gives an average repeatability score of 0.9582, surpassing other detectors by a significant margin (see Figure 1). The results for average keypoint density are presented in Table I, where DeepDetect achieves the highest average density of 0.5143. Table I also reports average foreground-to-keypoint ratio on $18$ custom-selected images with ground-truth binary masks. DeepDetect achieves good ratios, indicating that its keypoints are not only dense but also well-concentrated in semantically important regions in the images. With default thresholds, the classical detectors yield higher F-KP ratios but significantly lower keypoint densities. Conversely, reducing thresholds to increase density decreases the F-KP ratio, highlighting the presence of noisy keypoints. For instance, FAST and AGAST with low thresholds, achieve higher densities and lower F-KP ratios, showing their inability to focus salient regions. DeepDetect produces considerably higher densities while maintaining strong foreground focus. Visual comparisons in Figure 2 highlight how DeepDetect focuses on semantically meaningful regions and provides high keypoint density. Table II and Figure 6 show that DeepDetect provides substantially higher correct matches and consistently outperforms classical detectors. As far as learning based algorithms are concerned, Table III shows their repeatability comparison with DeepDetect on two different datasets. DeepDetect provides highest repeatabilities due to its intelligent keypoint-edge fusion pipeline. Similarly, Figure 7 shows its qualitative comparison with LIFT and SuperPoint on HPatches, highlighting that DeepDetect outperforms them with correct matches reaching as high as 338,118 for the busstop image pair. To further validate the benefits of DeepDetect, we compare its performance with SIFT (one of the best algorithms) for 3D reconstruction in Figure 8. It is evident that DeepDetect provides high quality, dense 3D reconstructions with number of voxels reaching as high as 842,045, substantially surpassing the performance of SIFT even when extremely low thresholds are used. The strength of DeepDetect lies in its ability to combine the geometric robustness and diversity of keypoint and edge detectors, and adaptability to extreme image variations gained through the use of supervision masks generated with “different thresholds” for normal and challenging scenes. DeepDetect demonstrates stronger generalization across various conditions without manual tuning due to its intelligent nature, making it well suited for visually volatile, changing, or degraded environments.